# Magnetic field and kinetic helicity evolution in simulations of interacting disk galaxies

### v1

@@ 5 Conclusions
We explored the evolution of galactic magnetic fields during interactions and mergers of disk galaxies by means of numerical simulations. Gaseous disks in hydrostatic equilibrium were initialized using the potential method of Wang et al. (2010). Moreover, we adopted the method of Drakos et al. (2017) to compute particle initial conditions for live dark matter halos. To compute the evolution of disks and halos, we employed the magnetohydrodynamical adaptive mesh refinement code Enzo. This allowed us to analyse magnetic field amplification induced by the tidal interactions between the galaxies. In an elaborate parameter study, we varied the impact parameter (i. e., the normal distance of the trajectory of the inbound galaxy to the target galaxy) and the relative orientations of disks.
@@
During interaction phases, we found pronounced peaks of the magnetic field strength averaged over center and outer regions of the disks. These peaks are a distinctive feature that was also identified by Drzazga et al. (2011) in radio observations of interacting galaxies. However, there are important differences. For all galaxies in their sample, Drzazga et al. (2011) found a relatively small increase during encounters, while the strongest fields are associated with the final coalescence phase. In our suite of simulations, this is only the case for impact parameters $\alpha_{\mathrm{b}}$ (the angle defined by the ratio of the linear impact parameter and the initial separation) greater than about $20^{\circ}$ or, to a lesser degree, for collisions of disks that are inclined with respect to their orbital motion and whose axes are not aligned (Fig. 6). Nearly central collisions typically exhibit the highest peaks during the first encounter. If the trends following from our simulations are correct, this implies that nearly central collisions are rare and were not present in the small galaxy sample investigated by Drzazga et al. (2011). They ordered the galaxies in a sequence of different interaction stages. However, our results indicates that, even at the same interaction stage, variations in the field strengths may result from different interaction parameters. Since three-dimensional data are required, it will be challenging to lift this degeneracy by classifying observed systems accordingly.
@@
Moreover, the mean field strength in center regions exceeds $10\;\upmu$G in about half of the scenarios we investigated, while the maximum mean field strength is typically in the range from $5$ to $10\;\upmu$G in off-center regions (Fig. 7). These values are systematically lower than the field strengths inferred by Drzazga et al. (2011) from non-thermal radio emission, which reach $50\;\upmu$G and $20\;\upmu$G in center and off-center regions, respectively. This discrepancy might be a consequence of insufficient numerical resolution, resulting in a less effective turbulent dynamo. However, comparable field strengths were obtained by Rodenbeck & Schleicher (2016), despite the significantly lower resolution of their simulations (240240240\;240pc compared to 151515\;15pc in our simulations). The most likely explanation is the assumption of adiabatic gas dynamics. Since the gas cannot cool, adiabatic disks do not contract into thin disks in our simulations. Thus, they have larger scale heights and the initial magnetic field undergoes less compression (assuming flux conservation), resulting in an overall lower field strength. This is supported by relative amplification factors of about $2$ to $3$ (Fig. 6), which compares well to Drzazga et al. (2011). In addition, magnetic field amplification due to turbulence driven by feedback processes is neglected. Whittingham et al. (2021) indeed reproduce observed galactic field strengths in cosmological zoom-in simulations with non-adiabatic physics and stellar feedback.
@@
As a quantitative indicator, Whittingham et al. (2021) showed that the power spectra of the magnetic field energy in the central regions of interacting galaxies approximately agree with Kolmogorov scaling toward small length scales. This suggests that the field is turbulent (Beresnyak 2019). However, power spectra are of limited validity for complex structures such as interacting galaxies, where the intensity of turbulence undergoes substantial spatial variations and rapid changes in time. Moreover, there are localized compression effects, for example, shock compression of gas in collision zones. For this reason, we applied a box filter to decompose the velocity and magnetic field into a mean-field part and small-scale fluctuations.
@@
The amplification of the magnetic field during encounters and coalescence is reflected by an enhanced EMF, which subsides in the aftermath of the interaction (Fig. 8, top row). As a result, the amplification is transient. Basically, this agrees with observational data of Drzazga et al. (2011) and the numerical studies (Pakmor et al. 2014; Rodenbeck & Schleicher 2016; Whittingham et al. 2021). By comparing small-scale and mean-field components, we find that the turbulent, smale-scale EMF grows after the first interaction to a fraction of the order of $0.1$ of the mean-field contribution. However, when the galaxies approach each other and get close to the orbital pericenter, the ratio of small-scale to mean-field EMF decreases in many cases (Fig. 8, bottom row). These minima are typically associated with antisymmetric peaks of the turbulent kinetic helicity above and below the midplane (Fig. 9). Antisymmetric kinetic helicity was previously reported by Ntormousi et al. (2020) for isolated disk galaxies. As pointed out by Brandenburg & Subramanian (2005), inhomogenous and anisotropic flows typically have large helicity and act as large-scale dynamo. Such flows are produced by tidal forces and, consequently, the mean-field contribution should be relatively large during close encounters. The small-scale EMF becomes more significant in post-interaction phases in which the gas becomes increasingly turbulent and when the galaxies finally merge. The turbulent kinetic helicity switches from antisymmetry to an irregular regime and decays in the remnant. If the rotation axis of both galaxies are inclined, such as in face-on collisions, a significant net helicity develops during the first encounter. In these scenarios, the maximal amplification of the magnetic field is relative low (Fig. 7).
@@
Overall, this leads to the interpretation that the peaks in the average magnetic field are mainly due to tidally induced helical flows. This process is more efficient if the disk plane is oriented parallel to relative motion. The production of turbulence enhances the magnetic field, especially in post-interaction phases. However, the small-scale dynamo is not sustained for long after coalescence. A lasting effect is the unordered, random structure of magnetic field in the remnants, which are always ellipsoids. However, in some cases this shape might be an artifact of adiabatic gas dynamics. It is conceivable that radiative cooling allows for disk-like remnants, where a coherent field could be rejuvenated by the $\alpha$-$\Omega$ dynamo. Apart from the highly dynamic evolution of magnetic fields, star formation and, in turn, feedback is triggered by galaxy collisions (Moreno et al. 2021; Whittingham et al. 2021; Linden & Mihos 2022; Renaud et al. 2022). Moreover, the dependence of the small-scale dynamo on numerical resolution in galaxy simulations is a pending problem (Grete et al. 2019; Körtgen et al. 2019; Ntormousi et al. 2020; Whittingham et al. 2021; Liu et al. 2022; Wissing & Shen 2023). As shown by Schober et al. (2013), the small-scale dynamo first amplifies the magnetic field to saturation on the viscous scale. However, this scale is in the sub-parsec range, far below scales that can be resolved in galaxy simulations. Subsequently, the magnetic energy is shifted to larger scales until the magnetic field saturates on the forcing scale. In a numerical simulation, the smallest scale is set by numerical diffusion. Consequently, the turbulent EMF computed from our simulations initially operates on scales that are orders of magnitude larger than the physical amplification scales. Numerical diffusion also determines the effective resistive scale on which magnetic energy is dissipated. Consequently, growth and decay rates may very well change with resolution. Since we demonstrated that the small-scale EMF plays a significant role in the turbulent environment of interacting galaxies, improved modelling of the underlying dynamo and dissipation processes will be an important next step.

### v2

@@ 5 Conclusions
We explored the evolution of galactic magnetic fields during interactions and mergers of disk galaxies by means of numerical simulations. Gaseous disks in hydrostatic equilibrium were initialized using the potential method from Wang et al. (2010). Moreover, we adopted the method from Drakos et al. (2017) to compute particle initial conditions for live dark-matter halos. To compute the evolution of disks and halos, we employed the MHD AMR code Enzo. This allowed us to analyze magnetic field amplification induced by the tidal interactions between the galaxies. In an elaborate parameter study, we varied the impact parameter (i.e., the normal distance of the trajectory of the inbound galaxy to the target galaxy) and the relative orientations of disks.
@@
During interaction phases, we find pronounced peaks in the magnetic field strength averaged over the center and outer regions of the disks. These peaks are a distinctive feature that was also identified by Drzazga et al. (2011) in radio observations of interacting galaxies. However, there are important differences. For all the galaxies in their sample, Drzazga et al. (2011) found a relatively small increase during encounters, and the strongest fields they found are associated with the final coalescence phase. In our suite of simulations, this is only the case for impact parameters $\alpha_{\mathrm{b}}$ (the angle defined by the ratio of the linear impact parameter and the initial separation) greater than about $20^{\circ}$ or, to a lesser degree, for collisions of disks that are inclined with respect to their orbital motion and whose axes are not aligned (Fig. 6). Nearly central collisions typically exhibit the highest peaks during the first encounter. If these trends are correct, this implies that nearly central collisions are rare and were not present in the small galaxy sample investigated by Drzazga et al. (2011), who ordered the galaxies by interaction stages. However, our results indicate that, even at the same interaction stage, variations in the field strengths may result from different interaction parameters. Since three-dimensional data are required, it will be challenging to lift this degeneracy by classifying observed systems accordingly.
@@
Moreover, the mean field strength in center regions exceeds $10\;\upmu$G in about half of the scenarios we investigated, while the maximum mean field strength is typically in the range from $5$ to $10\;\upmu$G in off-center regions (Fig. 7). These values are systematically lower than the field strengths inferred by Drzazga et al. (2011) from nonthermal radio emission, which reach $50\;\upmu$G and $20\;\upmu$G in center and off-center regions, respectively. This discrepancy might be a consequence of insufficient numerical resolution resulting in a less effective turbulent dynamo. However, comparable field strengths were obtained by Rodenbeck & Schleicher (2016) despite the significantly lower resolution of their simulations (240240240\;240pc compared to 151515\;15pc in our simulations). The most likely explanation is the assumption of adiabatic gas dynamics. Since the gas cannot cool, adiabatic disks do not contract into thin disks in our simulations. Thus, they have larger scale heights and the initial magnetic field undergoes less compression (assuming flux conservation), resulting in an overall lower field strength. This is supported by relative amplification factors of about $2$ to $3$ (Fig. 6), which compares well to the Drzazga et al. (2011) results. In addition, we neglected magnetic field amplification due to turbulence driven by feedback processes. Whittingham et al. (2021) do reproduce observed galactic field strengths in cosmological zoom-in simulations with nonadiabatic physics and stellar feedback.
@@
As a quantitative indicator, Whittingham et al. (2021) show that the power spectra of the magnetic field energy in the central regions of interacting galaxies approximately agree with Kolmogorov scaling toward small length scales. This suggests that the field is turbulent (Beresnyak 2019). However, power spectra are of limited validity for complex structures such as interacting galaxies, where the intensity of turbulence undergoes substantial spatial variations and rapid changes in time. Moreover, there are localized compression effects, for example shock compression of gas in collision zones. For this reason, we applied a box filter to decompose the velocity and magnetic field into a mean-field part and small-scale fluctuations.
@@
The amplification of the magnetic field during encounters and coalescence is reflected by an enhanced EMF, which subsides in the aftermath of the interaction (Fig. 8, top row). As a result, the amplification is transient. Basically, this agrees with the observational data from Drzazga et al. (2011) and the numerical studies (Pakmor et al. 2014; Rodenbeck & Schleicher 2016; Whittingham et al. 2021). By comparing small-scale and mean-field components, we find that the turbulent, small-scale EMF grows after the first interaction to a fraction on the order of $0.1$ of the mean-field contribution. However, when the galaxies approach each other and get close to the orbital pericenter, the ratio of small-scale to mean-field EMF decreases in many cases (Fig. 8, bottom row). These minima are typically associated with antisymmetric peaks of the turbulent kinetic helicity above and below the midplane (Fig. 9). Antisymmetric kinetic helicity was previously reported by Ntormousi et al. (2020) for isolated disk galaxies. As pointed out by Brandenburg & Subramanian (2005), inhomogeneous and anisotropic flows typically have large helicities and act as large-scale dynamos. Such flows are produced by tidal forces, and, consequently, the mean-field contribution should be relatively large during close encounters. The small-scale EMF becomes more significant in post-interaction phases in which the gas becomes increasingly turbulent and the galaxies finally merge. The turbulent kinetic helicity switches from antisymmetry to an irregular regime and decays in the remnant. If the rotation axis of both galaxies are inclined, such as in face-on collisions, a significant net helicity develops during the first encounter. In these scenarios, the maximal amplification of the magnetic field is relatively low (Fig. 7).
@@
Overall, this led to our interpretation that the peaks in the average magnetic field are mainly due to tidally induced helical flows. This process is more efficient if the disk plane is oriented parallel to relative motion. The production of turbulence enhances the magnetic field, especially in post-interaction phases. However, the small-scale dynamo is not sustained for long after coalescence. A lasting effect is the unordered, random structure of the magnetic field in the remnants, which are always ellipsoids. However, in some cases this shape might be an artifact of adiabatic gas dynamics. It is conceivable that radiative cooling allows for disk-like remnants, where a coherent field could be rejuvenated by the $\alpha$-$\Omega$ dynamo. Apart from the highly dynamic evolution of magnetic fields, star formation and, in turn, feedback are triggered by galaxy collisions (Moreno et al. 2021; Whittingham et al. 2021; Linden & Mihos 2022; Renaud et al. 2022). Moreover, the dependence of the small-scale dynamo on numerical resolution in galaxy simulations is a pending problem (Grete et al. 2019; Körtgen et al. 2019; Ntormousi et al. 2020; Whittingham et al. 2021; Liu et al. 2022; Wissing & Shen 2023). As shown by Schober et al. (2013), the small-scale dynamo first amplifies the magnetic field to saturation on the viscous scale. However, this scale is in the sub-parsec range, far below scales that can be resolved in galaxy simulations. Subsequently, the magnetic energy is shifted to larger scales until it saturates on the forcing scale. In a numerical simulation, the smallest scale is set by numerical diffusion. Consequently, the turbulent EMF computed from our simulations initially operates on scales that are orders of magnitude larger than the physical amplification scales. Numerical diffusion also determines the effective resistive scale on which magnetic energy is dissipated. Consequently, growth and decay rates may very well change with resolution. Since we have demonstrated that the small-scale EMF plays a significant role in the turbulent environment of interacting galaxies, improved modeling of the underlying dynamo and dissipation processes will be an important next step.